ADNAN MUHAMMED

Grade 12,

An 8000-N car is traveling at 12 m/s along a horizontal road when the brakes are applied. The car skids to a stop in 4.0 s. How much kinetic energy does the car lose in this time?

An 8000-N car is traveling at 12 m/s along a horizontal road when the brakes are applied. The car skids to a stop in 4.0 s. How much kinetic energy does the car lose in this time?

Grade:12th pass

2 Answers

Vikas TU
3 years ago
Dear student 
W = mg 
K = mv^2 /2 = Wv^2 /2g 
K = 8000 *12^2 /(2*9.8)
= 59 Kj 
Hope this will help 
Good Luck 
Yashraj
3 years ago
final kinetic energy is 0
initial kinetic energy is 1/2mv2
initial kinetic energy is 59 kilojoule
hope this help